• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Jing,Yuan

    Log entry time 00:08:56 on August17,2003

    Entry number 111341

    keyword=swing shift summary

    16:07 acceptance on the empty target ,raster on;
    16:17 acceptance on the empty target ,raster off;
    16:26 move target to C12 multifoil ;
    17:01 change the current and P0 of septum:I=59.8Amps;P0=0.475Gev/c;
    17:10 Bean off,MCC begin to change the pass (hopely 4 hours to finish)
    17:16 do the cycle of Q2and Q3 for both arms;
    17:18 move target to Pol.3He to do the NMR(39.1%),EPR(38.5%);
    17:38 change the current and P0 of septum:I=293.7Amps;P0=2.279Gev/c;
    left arm :Po=2.170;
    17:52 Move target to empty target;
    18:35 MCC call back ,they have finished setting 2 passes of accelerator.we asked control access to take out the collimator,and put in sieve.
    18:41 Pager Eugene back to do Moller measurement;
    19:22 Harp scan with 4 uA;H180,V 190;
    20:28 finish adjusting the position of beam;move target to C multifoils to begin OPTICS;
    22:00 Eugene begin to do Moller measurement ;
    23:55 there is no alarm of dipole from left arm after changing the P0 and current.

    Run number Field Target Beam plate Events(M) Raster Beam Energy

    23638/3638 0 Empty out 1.2/1.1 on 0.461
    23639/3639 0 Empty out 1.6/1.8 off 0.461
    23640/3640 0 12Cfoil out 2.1/2.0 off 0.461(prescale in left arm is too small ,deadtime is so large ,So we end it and have a new run)
    23641/3641 0 12Cfoil out 1.1/2.4 off 0.461
    I
    23642/3642 0 C foils out 0.78/1.0 off 2.279 293.3
    23643/3643 0 C foils out 0.5/1.0 off 2.213 290.0
    23644/3644 0 C foils out 0.2/0.5 off 2.213 284.3